Marian Plunket owns her own business and is considering an investment. If she undertakes the investment, it will pay $5,920 at the end of each of the next 3 years. The opportunity requires an initial investment of $1,480 plus an additional investment at the end of the second year of $7,400. What is the NPV of this opportunity if the interest rate is 1.8% per year? Should Marian take it?
Answer #1

Present value of annuity=Annuity[1-(1+interest rate)^-time period]/rate

=5920[1-(1.018)^-3]/0.018

=5920*2.895154673

=$17139.32

Present value of outflows=1480+7400*Present value of discounting factor(rate%,time period)

=1480+7400/1.018^2

=$8620.62

NPV=Present value of inflows-Present value of outflows

=$17139.32-$8620.62

=$8518.70(Approx).

Hence since NPV is positive;investment must be undertaken.